# Diff-Control: A Stateful Diffusion-based Policy for Imitation Learning

### v2

@@ V-B Duck Scooping Evaluation
Robustness Evaluation: Further evaluation was conducted to assess the consistency of the Diff-Control policy in this task. One way of quantifying action consistency is to measure the execution gap, which is the distance between the tail and head of two consecutive execution windows. In Figure 8(left), the 3D distance of the execution gap is illustrated for each policy network, with Diff-Control displaying the smallest gap.
Figure 8: Left: Analysis of the execution gap (measured in cm) for diverse tasks during evaluation. Right: Assessment of the success rate for each policy when faced with perturbations in the duck scooping task.
Furthermore, a supplementary set of experiments was carried out for the duck scooping task under visual occlusion. As depicted in Figure 8, Image-BC and ModAttn only achieved a 0% success rate when faced with perturbation, while BC-Z LSTM succeeded in 1 out of 10 trials in this scenario. Despite the diffusion policy achieving a 50% success rate, Diff-Control demonstrated an 80% success rate, showcasing the benefits of its stateful characteristic.
V-C Drum Beat Evaluation
@@
VI Conclusion
This study introduces Diff-Control, a stateful action diffusion policy designed for consistent action generation. The study explores the integration of diffusion model with ControlNet for robot action generation, demonstrating how temporal consistency can be enforced to enhance robustness and success rates. The efficacy of the proposed policy network is validated across four diverse real-world tasks, each highlighting a unique characteristic of the Diff-Control policy. Furthermore, our findings underscore the robustness and effectiveness of Diff-Control in managing dynamic and stateful tasks while remaining resilient against perturbations.
Limitation: The key limitations that require future consideration are as follows: 1) Diff-Control relies on the assumption that expert demonstrations are optimal, potentially leading to convergence at suboptimal states with insufficient demonstrations; 2) Diff-Control involves fine-tuning on a base policy to acquire action representations, necessitating a two-stage training process. Future work can explore automating the two-stage training process and manipulating different input conditions to ControlNet for other applications, i.e., torque sensors for contact-rich manipulation.
Acknowledgment
This work has received partial support from the National Science Foundation under grants CNS-1932068, IIS-1749783.

### v3

@@ IV-A Duck Scooping Evaluation
Figure 7: Left: Analysis of the execution gap (measured in cm) for duck scooping tasks during evaluation. Right: Assessment of the success rate for each policy when faced with perturbations in the duck scooping task.
Robustness Evaluation: Further evaluation was conducted to assess the consistency of the Diff-Control policy in this task. One way of quantifying action consistency is to measure the execution gap, which is the distance between the tail and head of two consecutive execution windows. In Figure 7(left), the 3D distance of the execution gap is illustrated for each policy network, with Diff-Control displaying the smallest gap. Furthermore, a supplementary set of experiments was carried out for the duck scooping task under visual occlusion. As depicted in Figure 7, Image-BC and ModAttn only achieved a 0% success rate when faced with perturbation, while BC-Z LSTM succeeded in 1 out of 10 trials in this scenario. Despite the diffusion policy achieving a 50% success rate, Diff-Control demonstrated an 80% success rate, showcasing the benefits of its stateful characteristic.
IV-B Drum Beat Evaluation
@@
V Conclusion
This study introduces Diff-Control, a stateful action diffusion policy designed for consistent action generation. The study explores the integration of diffusion model with ControlNet for robot action generation, demonstrating how temporal consistency can be enforced to enhance robustness and success rates. Furthermore, our findings underscore the robustness and effectiveness of Diff-Control in managing dynamic and stateful tasks while remaining resilient against perturbations.